# Adapting Image-to-Video Diffusion Models for Large-Motion Frame Interpolation

### v1

@@ Abstract
The development of video generation models has advanced significantly in recent years. For video frame interpolation, we adopt a pre-trained large-scale image-to-video diffusion model. To enable this adaptation, we propose a conditional encoder, which serves as a simple yet effective trainable module. By leveraging the first and last frames, we extract spatial and temporal features and input them into the conditional encoder. The computed features of the conditional encoder guide the video diffusion model in generating keyframe-guided video sequences. Our method demonstrates superior performance on the Fréchet Video Distance (FVD) metric compared to previous deterministic approaches in handling large-motion cases, highlighting advancements in generative-based methodologies.
@@ I Introduction
Building on the success of ControlNet[12] in the text-to-image domain, which has demonstrated exceptional control capabilities, we extend this method to keyframe-guided video frame interpolation. Experimental results indicate that this adaptation of ControlNet achieves great performance, maintaining temporal and spatial consistency while addressing the complexities of interpolation tasks. Moreover, to enhance temporal consistency in the interpolated frames, we integrate bidirectional optical flow into our approach, ensuring smoother and more coherent transitions between frames. Additionally, we incorporate a keyframe attention mechanism within each transformer block, allowing the model to better capture spatial and temporal representations by emphasizing keyframe feature. These enhancements improve the quality of the interpolated frames, enabling the model to handle complex motion dynamics and achieve better performance in video frame interpolation tasks.
@@
    We propose a dual-branch architecture to capture spatial-temporal features and integrate a keyframe attention module into each transformer block, enhancing interpolation performance.
@@ II-B Diffusion Model for Image and Video Generation
Diffusion models have demonstrated remarkable capabilities in the text-to-image domain, achieving state-of-the-art performance among generative models[17],[18],[19],[20],[21]. Building on their success in image generation, recent research has increasingly focused on applying diffusion models to video generation[22],[23]. Some approaches involve training a plug-and-play motion module that integrates seamlessly with existing text-to-image diffusion models[24], while others extend pre-trained text-to-image generation models by adding temporal modules to handle video dynamics[25],[26]. Both methods have shown promising results in synthesizing temporally coherent and visually compelling video content. Our approach aims to enhance control by fine-tuning image-to-video diffusion models through keyframe-guided generation.
@@ III-A Preliminaries
III-B Conditional Encoder
We present a novel plug-and-play conditional encoder module designed specifically for video frame interpolation shown in Fig 1. Firstly, we extract the first and last frames from the video data $\{I_{1},I_{2},\dots,I_{N}\}$, resulting in $\{I_{1},I_{N}\}$. We design spatial and temporal branches to extract spatial-temporal features, aiming to enhance interpolation performance shown in Fig 2. Additionally, a $1\times 1$ convolutional layer is employed to integrate conditional image CLIP [27] features for further processing.
Spatial Branch: Zero images are inserted between the first and last frames to represent the unconditional phase, with a binary mask concatenated to distinguish phases. The binary mask is set as $m=1$ for the conditional phase in the first and last frames, and $m=0$ for the unconditional phase in intermediate frames [28]. These components are concatenated along the channel dimension. A pre-trained autoencoder is then utilized to encode the combined inputs, extracting spatial features.
Temporal Branch: A pre-trained optical flow extractor [29] is employed to compute bi-directional optical flow, which are subsequently processed through a convolutional network to produce temporal features.
We adopt a simple early channel attention fusion strategy to combine spatial and temporal features [30], as the fused features will be processed by a complex conditional encoder capable of effectively learning the required representations autonomously. Specifically, the fusion of features is defined by the following equation:
$F_{\text{fused}}=\sigma\left(\text{MLP}(F_{s},F_{t})\right)\cdot F_{s}+\left(1-\sigma\left(\text{MLP}(F_{s},F_{t})\right)\right)\cdot F_{t}$ (2)
where $\sigma$ represents the sigmoid activation function, MLP is a multilayer perceptron consisting of convolutional layers, and $F_{s}$ and $F_{t}$ denote the spatial and temporal features, respectively. This equation adaptively balances the contributions of spatial and temporal features in the fused representation.
Figure 3: In both the spatial transformer block and the temporal transformer block, a keyframe attention block is first integrated to enhance feature representation.
@@ III-B Conditional Encoder
Ours 18.90 0.4701 0.3165 25.70 185.8 Ours 21.74 0.7077 0.2809 42.66 279.7
The design of the conditional encoder is inspired by ControlNet[12]. It replicates the encoder part of the pre-trained Stable Video Diffusion U-Net, initializing the weights identically. The conditional encoder comprises a convolutional block, a spatial transformer, and a temporal transformer block. In the spatial and temporal transformer blocks, we introduce keyframe attention to enhance the representation of both spatial and temporal fused features. Subsequently, the attention layer receives $n$ inputs, $z=[z^{1},\ldots,z^{n}]$, where $i=1,\ldots,n$. Through a linear projection operation, key and value features are derived from the key frames $z_{1},z_{n}$, which are concatenated to form $K^{(1,n)}$ and $V^{(1,n)}$ respectively. For each query feature from frame $z_{i}$, the $KeyframeAttn(Q,K,V)$ mechanism, as defined by the following equation:
$\text{Softmax}\left(\frac{Q\cdot({K^{1,n}})^{\top}}{\sqrt{d_{k}}}\right)V^{1,n}$ (3)
Specifically, in the spatial transformer block, conditional frames are utilized as key-value pairs to propagate appearance correlations from the spatial features. In the temporal transformer block, cross-frame interactions are modeled using conditional frames as key-value pairs to propagate implicit temporal coherence across other features, as illustrated in Fig. 3.
@@ IV-C Qualitative Result
We present the qualitative results, as illustrated in Fig. 6 Notably, when the first and last frames involve non-linear transformations or ambiguous motion, traditional methods such as AMT and FILM fail to generate accurate intermediate frames. In contrast, our model generates superior results, characterized by accurate motion and clear textures. Additionally, our method outperforms SVDKFI in low-resolution scenarios, demonstrating greater robustness to the effects of down sampling. Furthermore, our approach exhibits the capability to interpolate across diverse image styles. We evaluate its performance on real-world, anime, and sketch styles, as illustrated in Fig. 4.
Figure 5: The example illustrates a large-motion scenario. Whereas other methods struggle to accurately predict motion, our approach successfully reconstructs and captures realistic motion dynamics.
IV-D Ablation study
We conducted experiments on the proposed components of our method to perform an ablation study. Refer to Table II for detailed results. For a fair comparison of the proposed method, we discard the $256\times 256$ versions of the DAVIS-7 and UCF101-7 datasets. Instead, we re-crop the DAVIS 2017 dataset to $512\times 512$ resolution, retaining sequences of 7 frames, where the model is tasked with predicting the middle 5 frames. The baseline model includes only the spatial branch, enabling an image-to-video model to adapt to video frame interpolation. In contrast, our method integrates the temporal branch and the keyframe attention module. Experimental results demonstrate that the proposed components enhance the performance of the baseline model.
Figure 6: Effect of resolution on frame interpolation performance. The left example depicts a middle frame at $512\times 512$ resolution, while the right example is at $256\times 256$ resolution.
@@ IV-D Ablation study
Ours 19.74 0.5214 0.3130 16.32 122.6
IV-E Effect of Resolution on Interpolation Quality
In this section, we analyze the impact of resolution on the generation quality of Stable Video Diffusion. Stable Video Diffusion is a type of latent diffusion model that employs an autoencoder to down sample data into a latent space, facilitating more efficient training. However, this down sampling process is inherently lossy and can lead to suboptimal performance when generating low-resolution images. As illustrated in Fig. 6, the results highlight the influence of different resolutions. In particular, the first and last frames are utilized as conditional frames, but the down sampling process results in the loss of pixel-level information, thereby impairing the accurate prediction of the middle frame.
@@ VI Conclusion
We introduce a conditional encoder designed to adapt an image-to-video model for large-motion frame interpolation. To enhance performance, we integrate a dual-branch feature extractor and propose a keyframe attention mechanism that effectively captures both spatial and temporal information, enabling accurate interpolation of intermediate frames. Despite its simplicity, our method proves highly effective, achieving comparable results on the Fréchet Video Distance (FVD) metric when evaluated against other state-of-the-art approaches.

### v2

@@ Abstract
With the development of video generation models has advanced significantly in recent years, we adopt large-scale image-to-video diffusion models for video frame interpolation. We present a conditional encoder designed to adapt an image-to-video model for large-motion frame interpolation. To enhance performance, we integrate a dual-branch feature extractor and propose a cross-frame attention mechanism that effectively captures both spatial and temporal information, enabling accurate interpolations of intermediate frames. Our approach demonstrates superior performance on the Fréchet Video Distance (FVD) metric when evaluated against other state-of-the-art approaches, particularly in handling large motion scenarios, highlighting advancements in generative-based methodologies.
@@ I Introduction
Building on the success of ControlNet[12] in the text-to-image domain, which has demonstrated exceptional control capabilities, we extend this method to keyframe-guided video frame interpolation. Experimental results indicate that this adaptation of ControlNet achieves great performance, maintaining temporal and spatial consistency while addressing the complexities of interpolation tasks. Moreover, to enhance temporal consistency in the interpolated frames, we integrate bidirectional optical flow and depth map into our approach, ensuring smoother and more coherent transitions between frames. Additionally, we incorporate a cross-frame attention mechanism within each transformer block, allowing the model to better capture spatial and temporal representations by emphasizing cross-frame feature. These enhancements improve the quality of the interpolated frames, enabling the model to handle complex motion dynamics and achieve better performance in video frame interpolation tasks.
@@
    We propose a dual-branch architecture to capture spatial-temporal features and integrate a cross-frame attention module into each transformer block, enhancing interpolation performance.
@@ II-B Diffusion Model for Image and Video Generation
Diffusion models are a type of generative model that have demonstrated remarkable capabilities in the text-to-image domain, achieving state-of-the-art performance among generative approaches [17],[18],[19],[20],[21]. Building on their success in image generation, recent research has increasingly focused on applying diffusion models to video generation[22],[23]. Some approaches involve training a plug-and-play motion module that integrates seamlessly with existing text-to-image diffusion models[24], while others extend pre-trained text-to-image generation models by adding temporal modules to handle video dynamics[25],[26]. Both methods have shown promising results in synthesizing temporally coherent and visually compelling video content.
@@ III-A Preliminaries
Figure 3: In both the spatial transformer block and the temporal transformer block, a cross-frame attention block is first integrated to enhance feature representation.
@@
Ours 19.53 0.4915 0.3064 24.84 177.79 Ours 23.09 0.7541 0.2501 38.68 265.74
III-B Conditional Encoder
We present a novel plug-and-play conditional encoder module designed specifically for video frame interpolation shown in Fig 1. Firstly, we extract the first and last frames from the video data $\{I_{1},I_{2},\dots,I_{N}\}$, which results in $\{I_{1},I_{N}\}$. We design spatial and temporal branches to extract spatial-temporal features, aiming to enhance interpolation performance shown in Fig. 2. Additionally, a $1\times 1$ convolutional layer is employed to integrate conditional image CLIP [27] features for further processing.
Spatial Branch: Zero images are inserted between the first and last frames to represent the unconditional phase, with a concatenated binary mask to distinguish phases. The binary mask is set as $m=1$ for the conditional phase in the first and last frames, and $m=0$ for the unconditional phase in intermediate frames [28]. These components are concatenated along the channel dimension. A pre-trained autoencoder is then utilized to encode the combined inputs, extracting spatial features.
Temporal Branch: A pre-trained optical flow extractor [29] and a depth extractor [30] are utilized to compute bi-directional optical flow and the keyframe depth map. These motion representations are subsequently processed through a convolutional network to produce temporal features. The resulting temporal features indicate keyframe occlusions, aiding in the interpolation modeling process.
We adopt a simple early fusion strategy to combine spatial and temporal features, as the fused features will be processed by a complex conditional encoder capable of effectively learning the required representations autonomously. Specifically, the fusion of features is defined by the following equation:
$\mathbf{F_{s}}=\text{CBAM}(\mathbf{F_{s}}),\quad\mathbf{F_{t}}=\text{CBAM}(\mathbf{F_{t}}),$ (2)
$\mathbf{F_{used}}=\text{CrossAttention}(\mathbf{F_{s}}\cdot\mathbf{W}_{Q},\mathbf{F_{t}}\cdot\mathbf{W}_{K},\mathbf{F_{t}}\cdot\mathbf{W}_{V})$ (3)
$\mathbf{F}_{\text{used}}=\mathbf{F}_{\text{used}}+\mathbf{F}_{\text{s}}$ (4)
Where $F_{s}$ and $F_{t}$ denote the spatial and temporal features, respectively. The spatial and temporal features $F_{s}$ and $F_{t}$ are passed through the CBAM[31] to obtain corresponding weights, which are then applied adaptively to the spatial-temporal features. These combined features are fed into the Cross Attention for feature fusion. This approach adaptively balances the contributions of spatial and temporal features in the fused representation.
The design of the conditional encoder is inspired by ControlNet[12]. It replicates the encoder part of the pre-trained Stable Video Diffusion U-Net, initializing the weights identically. The conditional encoder comprises a convolutional block, a spatial transformer, and a temporal transformer block. In the spatial and temporal transformer blocks, we introduce cross-frame attention to enhance the representation of both spatial and temporal fused features. The attention layer processes $n$ inputs, denoted as $z=[z^{1},\ldots,z^{n}]$, where $i=1,\ldots,n$. Linear projection operations are applied to generate query, key, and value features. In the spatial transformer block, for each query feature derived from a frame $z_{i}$, we utilize the first and last frames as keyframes, denoted as $z_{1}$ and $z_{n}$. These features are concatenated to form $K^{(1,n)}$ and $V^{(1,n)}$, respectively. The cross-frame attention mechanism, $CrossFrameAttn(Q,K,V)$, is mathematically defined as follows.
$\text{Softmax}\left(\frac{Q\cdot({K^{(1,n)}})^{\top}}{\sqrt{d_{k}}}\right)V^{(1,n)}$ (5)
Similarly, in the temporal transformer block, for each query feature derived from a frame $z_{i}$, as well as the preceding and succeeding frames, $z_{i-1}$ and $z_{i+1}$, these features are concatenated to form $K^{(i-1,i+1)}$ and $V^{(i-1,i+1)}$, respectively. The cross-frame attention mechanism, $CrossFrameAttn(Q,K,V)$, is mathematically defined as follows.
$\text{Softmax}\left(\frac{Q\cdot({K^{(i-1,i+1)}})^{\top}}{\sqrt{d_{k}}}\right)V^{(i-1,i+1)}$ (6)
Specifically, in the spatial transformer block, conditional frames are utilized as key-value pairs to propagate keyframe appearance correlations from the spatial features. In the temporal transformer block, cross-frame interactions are modeled using conditional frames as key-value pairs to propagate implicit temporal coherence across other features, as illustrated in Fig. 3.
@@ IV-A Experiment Settings
Figure 5: The example demonstrates a large-motion scenario where traditional methods fail to predict the direction and magnitude of motion accurately, often resulting in blurred or inconsistent intermediate frames. In contrast, our approach reconstruct realistic motion trajectories, producing coherent intermediate frames that capture accurate motion dynamics.
@@ IV-C Qualitative Result
We present the qualitative results, as illustrated in Fig. 5. Notably, when the first and last frames involve non-linear transformations or large motion, traditional methods such as AMT and FILM fail to generate accurate intermediate frames. In contrast, our model generates superior results, characterized by accurate motion and clear textures. Furthermore, our method outperforms SVDKFI in low-resolution scenarios, demonstrating greater robustness to the effects of down sampling. Furthermore, our approach exhibits the capability to interpolate across diverse image styles. We evaluate its performance in real-world, anime, and sketch styles, as illustrated in Fig. 4.
@@
Ours 20.28 0.5375 0.3069 16.27 118.05
IV-D Ablation study
We conducted experiments on the proposed components of our method to perform an ablation study. Refer to Table II for detailed results. For a fair comparison of the proposed method, we discard the $256\times 256$ versions of the DAVIS-7 and UCF101-7 datasets. Instead, we re-crop the DAVIS 2017 dataset to $512\times 512$ resolution, retaining sequences of 7 frames, where the model is tasked with predicting the middle 5 frames. The baseline model includes only the spatial branch, enabling an image-to-video model to adapt to video frame interpolation. In contrast, our method integrates the temporal branch and the cross-frame attention module. Experimental results demonstrate that the proposed components enhance the performance of the baseline model.
@@ VI Conclusion
We introduce a conditional encoder designed to adapt an image-to-video model for large-motion frame interpolation. To enhance performance, we integrate a dual-branch feature extractor and propose a cross-frame attention mechanism that effectively captures both spatial and temporal information, enabling accurate interpolation of intermediate frames. Despite its simplicity, our method proves highly effective, achieving comparable results on the Fréchet Video Distance (FVD) metric when evaluated against other state-of-the-art approaches.